OpenStudy (anonymous):

Find the exact value of the expression. tan(cos^-1(root3/2))

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

is this your question? ? : \[\huge{tan(cos^{-1}(\sqrt{\frac{3}{2}}))}\]

mathslover (mathslover):

was that your question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

cos^-1 represents inverse or just 1/cos ?

OpenStudy (anonymous):

inverse

mathslover (mathslover):

so let : \[\huge{tan(cos^{-1}(\sqrt{\frac{3}{2}}))=tan(x) }\] where x = cos^{-1}(root(3/2))

mathslover (mathslover):

ok till now?

OpenStudy (anonymous):

yeah

mathslover (mathslover):

so as we know that cos x = -\(\large{\frac{\sqrt{3}}{2}}\) cos^2x = ? can u tell me

OpenStudy (anonymous):

what is the need of cos^2x? @mathslover

mathslover (mathslover):

there is the need you just tell me what will be cos^2 x if cos x = - root3 /2

mathslover (mathslover):

u there @KingAditya ?

OpenStudy (anonymous):

yeah

mathslover (mathslover):

answer my question aditya... what will be cos^2 x if cos x = - root3 /2

OpenStudy (anonymous):

wait

mathslover (mathslover):

ok

mathslover (mathslover):

@KingAditya r u confused? tell me : \[\large{\textbf{what is :}(\frac{-\sqrt{3}}{2})^2}\]

OpenStudy (anonymous):

i think it might be 3/4

OpenStudy (anonymous):

cos^2=3/4

OpenStudy (anonymous):

cos^2x=3/4

mathslover (mathslover):

oh great .. now cos^2 x = 1 - sin^2 x so : 3/4 = 1 - sin^2 x 1 - 3/4 = sin^2 x 1/4 = sin^2 x sin x = 1/2 right?

mathslover (mathslover):

\[\large{tan x = \frac{sin x}{cos x}}\] \[\large{tan x = \frac{-1}{\sqrt{3}}}\]

mathslover (mathslover):

that is ur answer.. got it?

OpenStudy (anonymous):

1/sqrt(3)??

mathslover (mathslover):

-1/sqrt{3} sorry for that mistake

OpenStudy (anonymous):

-??

mathslover (mathslover):

yes -1/sqrt{3}

mathslover (mathslover):

it might be +ve if theta is in 3rd quadrant

OpenStudy (anonymous):

u mean in T quadrant??

mathslover (mathslover):

tan is +ve in third quadrant

mathslover (mathslover):

got it?

OpenStudy (anonymous):

@mathslover , he said cos^-1 represents the inverse, so it is arccos root3/2 which is Pi/6 . And tan(Pi/6) is 1/root3 = root3/3. That's it. @KingAditya

mathslover (mathslover):

oh yes right myko